OpenStudy (anonymous):

If the midpoint between (x, 7) and (10, 11) is (15, 9), find the value of x.

OpenStudy (anonymous):

will i have to use a square root?

ganeshie8 (ganeshie8):

midpoint = \(\large (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2}) \)

ganeshie8 (ganeshie8):

no, this is simpler than that... (x, 7) and (10, 11) x1 y1 x2 y2

ganeshie8 (ganeshie8):

x coordinate if midpoint : \(\large \frac{x+10}{2} = 15 \)

ganeshie8 (ganeshie8):

you can solve x

ganeshie8 (ganeshie8):

??

OpenStudy (anonymous):

umm i dont know how the 15 is there

ganeshie8 (ganeshie8):

15 is the midpoint's x-coordinate. we're given that...

ganeshie8 (ganeshie8):

\(\large \frac{x+10}{2} = 15\) multiply 2 both sides \(\large x+10 = 30\)

ganeshie8 (ganeshie8):

rest il leave it to you :)

OpenStudy (anonymous):

30-10=20

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

20/2=10 and then thats my coordinate or no?

ganeshie8 (ganeshie8):

x = 20 is ur answer

OpenStudy (anonymous):

okay thankss lovee bugg!

ganeshie8 (ganeshie8):

np :))